OpenStudy (ray1998):

Use graphing to find the solutions to the system of equations. x^2 - y = 4 2x + y = -1 Will give fan and medal. Thank you in advance!!

OpenStudy (ray1998):

Hey @Alphabet_Sam , can you show me how to start this?

OpenStudy (campbell_st):

what graphing tools do you have..?

OpenStudy (anonymous):

Sure. Do you understand how to graph either, or do you need help with both?

OpenStudy (ray1998):

Both please :)

OpenStudy (campbell_st):

just graph the 2 curves here then read off the point(s) of intersection https://www.desmos.com/calculator

OpenStudy (anonymous):

Are you allowed to use a graphing calculator?

OpenStudy (ray1998):

Graphing paper and a pencil is what they normally have you use @campbell_st

OpenStudy (ray1998):

Oh! Yes. Although, I don't have one with me

OpenStudy (campbell_st):

ok...so is this algebra or calculus..?

OpenStudy (ray1998):

Algebra 2 @campbell_st

OpenStudy (anonymous):

What @campbell_st linked is a graphing calculator. You can just plug the two formulas in there, and then the solution is where the two graphs touch.

OpenStudy (campbell_st):

ok... so you need to solve the quadratic equation to find where it cuts the x-axis I have rewritten the equation. Can you solve \[y = x^2 - 4\] by setting y = 0

OpenStudy (ray1998):

Okay :) thank you. @Alphabet_Sam

OpenStudy (ray1998):

I think so, yes. @campbell_st

OpenStudy (campbell_st):

ok... so you need to solve \[0 = x^2 - 4\] it can be factored

OpenStudy (anonymous):

Before you do that, the question asks you to solve by graphing. That's solving algebraically. It will produce the same answer, but if you are graded on your method it would be wrong.

OpenStudy (campbell_st):

to graph the curve you need to find some key information.... such as zeros or intercepts...

OpenStudy (anonymous):

Oh, you're right you can do it that way.

OpenStudy (ray1998):

Is this the correct answer?

OpenStudy (ray1998):

Where the points intersect from the formulas

OpenStudy (ray1998):

Okay. Thank you guys @Alphabet_Sam and @campbell_st :) you guys are great.

OpenStudy (campbell_st):

is that a guess or you worked it out..?

OpenStudy (ray1998):

I used the graphing calculator link that you posted.

OpenStudy (triciaal):

looks right

OpenStudy (campbell_st):

lol.... ok.... so much for the pencil and paper idea.... and its correct

OpenStudy (ray1998):

Lol normally they have you do that or work on a calculator :) so either way. Thank you so much!